# Compliance-to-Code: Enhancing Financial Compliance Checking via Code Generation

### v2

@@ Abstract
Nowadays, regulatory compliance has become a cornerstone of corporate governance, ensuring adherence to systematic legal frameworks. At its core, financial regulations often comprise highly intricate provisions, layered logical structures, and numerous exceptions, which inevitably result in labor-intensive or comprehension challenges. To mitigate this, recent Regulatory Technology (RegTech) and Large Language Models (LLMs) have gained significant attention in automating the conversion of regulatory text into executable compliance logic. However, their performance remains suboptimal particularly when applied to Chinese-language financial regulations, due to three key limitations: (1) incomplete domain-specific knowledge representation, (2) insufficient hierarchical reasoning capabilities, and (3) failure to maintain temporal and logical coherence. One promising solution is to develop a domain specific and code-oriented dataset for model training. Existing datasets such as LexGLUE, LegalBench, and CODE-ACCORD are often English-focused, domain-mismatched, or lack fine-grained granularity for compliance code generation. To fill these gaps, we present Compliance-to-Code  [^1] https://github.com/AlexJJJChen/Compliance-to-Code, the first large-scale Chinese dataset dedicated to financial regulatory compliance. Covering 1,159 annotated clauses from 361 regulations across ten categories, each clause is modularly structured with four logical elements-subject, condition, constraint, and contextual information-along with regulation relations. We provide deterministic Python code mappings, detailed code reasoning, and code explanations to facilitate automated auditing. To demonstrate utility, we present FinCheck: a pipeline for regulation structuring, code generation, and report generation. Compliance-to-Code establishes a new benchmark for LLM-based compliance automation. Experimental evaluation shows that GLM-4-9B-0414 achieves the best performance on key task in regulation structuring and Qwen3-8B performs the best in compliance code generation tasks.
footnotetext: *These authors contributed equally to this work.footnotetext: †Corresponding Author
@@ 1 Introduction
Corporate compliance with financial regulations is regarded as a critical mechanism through which institutions ensure their operations adhere to complex legal frameworks and industry standards (DePamphilis, 2019). Modern financial regulations are notoriously elaborate, featuring multilayered conditions, exceptions, quantitative thresholds, and numerous cross-references, all of which require accurate interpretation and operationalization (Butler and O’Brien, 2019). Traditionally, compliance checking relies heavily on legal, risk, and audit teams, incurring high costs and lengthy processing times. As the volume and complexity of regulations continue to grow, these manual approaches have become increasingly unsustainable (Weirich et al., 2025; Martin et al., 2024).
Regulatory Technology (RegTech) aims to automate compliance management by leveraging advanced computational techniques (Charoenwong et al., 2024). With the emergence of large language models (LLMs), substantial progress has been made in various legal text processing tasks, such as information retrieval, classification, and summarization (Guha et al., 2023; Kant et al., 2025). However, directly employing LLMs for compliance checking poses significant challenges. Due to the inherently black-box nature of LLMs and the non-determinism of their outputs, the generated answers may lack consistency and transparency. By contrast, executable code offers clear, verifiable compliance decisions. This raises a fundamental question: Can LLMs bridge the gap between natural language regulations and executable compliance code, thereby enabling fully automated compliance auditing? To explore this, we conducted preliminary experiments with state-of-the-art LLMs, including Gemini 2.5 Pro (Kavukcuoglu, 2025) and Qwen3-235B (Qwen Team, 2025). As illustrated in Figure 1, our results indicate that existing LLMs still exhibit significant limitations in regulation-to-code tasks:
  • Incomplete Domain Knowledge: LLMs may conflate domain concepts or overlook essential calculation steps. For instance, Gemini 2.5 Pro detected stock distributions by monitoring overall capital changes, ignoring whether such changes resulted strictly from bonus shares, and miscalculated diluted EPS by skipping post-distribution adjustments, resulting in flawed compliance judgments.
@@
  • Temporal and Logical Inconsistencies: Both models demonstrate limitations in tracking event causality and enforcing regulatory constraints over time. Qwen3-235B defaulted to over-simplified constraint satisfactions in the absence of explicit disclosure data, while Gemini 2.5 Pro wrongly applied event flags across all future periods and relied on ambiguous keyword searches in announcements, failing to uphold event-specific, rule-based compliance.
We attribute these weaknesses primarily to the insufficient domain-specific knowledge of LLMs, and inability to modularize regulatory content, and the absence of appropriate code-oriented training data. Previous research has attempted to address these issues by constructing datasets for improved LLM evaluation in compliance code generation. Existing benchmarks-such as LexGLUE (Chalkidis et al., 2022), LegalBench (Guha et al., 2023), and MAUD (Wang et al., 2023)-focus on tasks like legal text classification, clause extraction, or high-level reasoning. However, they do not modularize financial regulations, and therefore are inadequate for the end-to-end compliance code generation scenario (Küçük and Can, 2025). Other datasets, such as CODE-ACCORD (Hettiarachchi et al., 2025), attempt to formalize condition-action relationships within regulations, but are largely domain-agnostic and not tailored to financial compliance, which frequently relies on specialized terminology. Furthermore, CODE-ACCORD (Hettiarachchi et al., 2025) and similar datasets do not effectively reduce model hallucinations in this domain and are predominantly in English, limiting their applicability to diverse regulatory environments. Consequently, there is a lack of a high-quality, Chinese-language dataset for financial compliance checking-a critical resource for advancing LLMs’ understanding, reasoning, and automation capabilities in regulatory contexts.
@@
Experimental Results: We benchmark recent LLMs for regulation structuring and code generation tasks using our dataset’s detailed annotations. Qwen3-8B (Yang et al., 2025) achieves the best performance on code generation. With Supervised Fine-Tuning (SFT), GLM-4-9B-0414 notably improves its regulation structuring capabilities.
@@ 2.1 RegTech Methods for Financial Compliance
Regulatory Technology (RegTech) leverages computational tools to automate financial compliance (Butler and O’Brien, 2019). Early RegTech solutions relied on deterministic, rule-based systems and expert models (e.g., GEM (Hassan and Logrippo, 2009)) using formal logic and standards like XBRL (Jeyasingh, 2023). While precise and explainable, these symbolic systems struggled with scalability and adaptability amid evolving regulations.
Recent advances in AI, notably Large Language Models (LLMs), have enabled data-driven approaches for parsing regulations and automating compliance workflows (Sun et al., 2025a; Jiang et al., 2025). These neural models offer broader generalization but face challenges with domain specificity, rapid regulatory change, and the complexity of legal logic. Key limitations remain in transparency, interpretability, and auditability-essential for compliance contexts (Wang et al., 2024a).
To address this, hybrid pipelines combining neural models with symbolic reasoning or human oversight have become prominent (Wang et al., 2024a; Sun et al., 2025b). Our work follows this trajectory, presenting a code-centric approach for LLM-driven compliance automation that enhances rigor, reliability, and legal traceability beyond brittle rule-based pipelines.
@@ 2.2 Related Datasets and Benchmarks
Despite active research in RegTech, there is a critical lack of datasets mapping real regulatory text to executable compliance logic in the financial sector. Most existing resources address subtasks such as document classification (e.g., LEDGAR (Tuggener et al., 2020), LexGLUE (Chalkidis et al., 2022)), judgment prediction, or QA (e.g., LegalBench (Guha et al., 2023), BillSum (Kornilova and Eidelman, 2019)), but lack annotations for code-level reasoning or logic synthesis.
Datasets closer to financial regulation, like ObliQA (Gokhan et al., 2024) and BSARD (Lotfi et al., 2025), only support question answering or retrieval. CODE-ACCORD (Hettiarachchi et al., 2025) stands out as a logic-annotated resource, but it is limited to building regulations and has narrow domain coverage. Crucially, no existing benchmark, especially in Chinese language systematically couples financial statutes with executable, expert-annotated compliance logic, impeding progress towards true compliance to code. Our work directly addresses this foundational gap.
@@ 3.1 Data Collection and Pre-processing
Compliance-to-Code Dataset is systematically curated from official regulatory documents published by the Beijing Stock Exchange (BSE) [^2] https://www.bse.cn/rule/secnotice_list.html, with a focus on compliance obligations, internal controls, and risk points for publicly listed companies in China. In comparison to the regulations issued by the China Securities Regulatory Commission (CSRC) [^3] http://www.csrc.gov.cn, the BSE provides more detailed and supplementary rules, increasing both the granularity and complexity of regulatory interpretation. We specifically gathered authoritative documents classified as Departmental Rules and Normative Documents, which constitute the regulatory backbone of corporate compliance obligations and reporting requirements. The regulatory corpus reflects all on-force documents as of January 1, 2025, providing a stable but comprehensive snapshot for benchmarking compliance automation.
@@ 3.2 Data Annotation
Drawing inspiration from RegGPT (Wang et al., 2024b), we observe that regulations are typically composed of six core elements: meta data (entity/property), constraint, condition, measure, scope, and external references. To facilitate more effective translation of regulatory rules into executable code by LLMs, we propose several refinements to this schema. Specifically, we merge the elements of measure and constraint under the unified category of Constraint, relabel external references as Contextual Information, and rename meta data to Subject.
@@
We reformulated each programming problem as a code completion task and categorized them into three difficulty levels: simple, medium, and difficult. For Simple problems, the LLM needed to fill in one or two missing lines of code; for Medium problems, three lines; and for Difficult problems, the LLM was tasked with generating the entire solution from scratch. DeepSeek-R1 (Guo et al., 2025) provided candidate extractions and code suggestions, but all outputs underwent independent expert review and adjudication.
@@ 3.3 Dataset Statistics
Final annotation produced 1,159 CUs (from 1,226 candidates), and 864 directed inter-unit relations - refer to (61.1%), exclude (2.1%), only include (19.7%), should include (17.1%). Each CU includes averages 0.72 relations, 0.95 subjects, 0.80 conditions, and 0.92 constraints and 0.18 contextual informations. (in Figure 2)
These were mapped to 307 distinct executable compliance automation tasks (Python), each code module has COT reasoning step (average 8 steps/2145 tokens). Tasks stratify by scenario complexity: Simple (70.03%), Medium (11.07%), and Difficult (18.89%).
@@ 4 FinCheck: A Structured Compliance Checking Pipeline
Figure 3: FinCheck Compliance Checking Pipeline. The framework processes natural language regulations by first using a Structure Predictor to extract key compliance units. These units are then fed into a Code Generator to create verification code. For a specific case, user input triggers an Information Retriever to fetch relevant company data. The generated code is executed with this data, and a Report Generator summarizes the outcome before the final verification result is shown to user.
@@ 5.2 Baselines
Since there are currently no models specifically designed to convert regulations into code, we are unable to employ any models directly aimed at Task 2. Furthermore, although RegGPT (Wang et al., 2024b) demonstrates promising performance for Task 1, it is not open-sourced, which limits our ability to use it in our experiments. As a result, we focus our experiments on general LLMs.
@@ 5.3 Dataset Splits
Our dataset comprises 1,159 compliance units (CUs) and 307 compliance automation tasks. We divide the data into training (70%), validation (15%), and test (15%) sets, ensuring stratification by thematic area and task complexity (Simple, Medium, and Difficult). For model training, we exclusively utilize tasks labeled as Difficult, while evaluation is conducted on all three complexity levels within the test set.
@@ 5.4 Training Protocols
Base Models: For Task 1, we utilize Qwen3-8B (Qwen Team, 2025), DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), GLM-4-9B-0414 (THUDM, 2025), GPT-4.1 (OpenAI, 2025), Gemini-pro 2.5 (Kavukcuoglu, 2025), and Claude-3.7 (Anthropic, 2025). For Tasks 2 and 3, which focus on code generation, we additionally include Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) and DeepSeek-Coder-6.7B-Instruct (Guo et al., 2024) in our evaluation.
Supervised Fine-Tuning (SFT): We train using causal language modeling (next-token prediction), AdamW optimizer, cosine learning rate schedule, batch size 1, weight decay 0.01. The learning rate is set at $1\times 10^{-4}$, epochs $1$, top_p $0.75$, and temperature 00. For LoRA, we use rank 8, $\alpha=32$, and dropout $0.05$. We employ Swift  [^5] https://github.com/modelscope/ms-swift for efficiency. The prompt is shown in Appendix H.
@@ 5.5 Evaluation Metrics
For Tasks 2 & 3 (Code Generation), we primarily evaluate Functional Correctness, measured by Pass@1 and Pass@5. Specifically, Pass@k represents the percentage of generated code samples that successfully compile and pass all associated unit tests within the first k attempts. The temperature was set to 0.6 for decoding. As a secondary metric, we report Code Similarity, quantified by CodeBLEU (Ren et al., 2020), which comprehensively assesses the lexical, syntactic, and semantic similarity between generated code and the reference solutions.

### v3

@@ Abstract
Nowadays, regulatory compliance has become a cornerstone of corporate governance, ensuring adherence to systematic legal frameworks. At its core, financial regulations often comprise highly intricate provisions, layered logical structures, and numerous exceptions, which inevitably result in labor-intensive or comprehension challenges. To mitigate this, recent Regulatory Technology (RegTech) and Large Language Models (LLMs) have gained significant attention in automating the conversion of regulatory text into executable compliance logic. However, their performance remains suboptimal particularly when applied to Chinese-language financial regulations, due to three key limitations: (1) incomplete domain-specific knowledge representation, (2) insufficient hierarchical reasoning capabilities, and (3) failure to maintain temporal and logical coherence. One promising solution is to develop a domain specific and code-oriented dataset for model training. Existing datasets such as LexGLUE, LegalBench, and CODE-ACCORD are often English-focused, domain-mismatched, or lack fine-grained granularity for compliance code generation. To fill these gaps, we present Compliance-to-Code  [^1] https://github.com/AlexJJJChen/Compliance-to-Code, the first large-scale Chinese dataset dedicated to financial regulatory compliance. Covering 1,159 annotated clauses from 361 regulations across ten categories, each clause is modularly structured with four logical elements-subject, condition, constraint, and contextual information-along with regulation relations. We provide deterministic Python code mappings, detailed code reasoning, and code explanations to facilitate automated auditing. To demonstrate utility, we present FinCheck: a pipeline for regulation structuring, code generation, and report generation. Compliance-to-Code establishes a new benchmark for LLM-based compliance automation. Experimental evaluation shows that GLM-4-9B-0414 achieves the best performance on key task in regulation structuring and DeepSeek-R1-0528 performs the best in compliance code generation tasks.
@@ 1 Introduction
Corporate compliance with financial regulations is regarded as a critical mechanism through which institutions ensure their operations adhere to complex legal frameworks and industry standards DePamphilis (2019). Modern financial regulations are notoriously elaborate, featuring multilayered conditions, exceptions, quantitative thresholds, and numerous cross-references, all of which require accurate interpretation and operationalization Butler and O’Brien (2019). Traditionally, compliance checking relies heavily on legal, risk, and audit teams, incurring high costs and lengthy processing times. As the volume and complexity of regulations continue to grow, these manual approaches have become increasingly unsustainable Weirich et al. (2025); Martin et al. (2024).
Regulatory Technology (RegTech) aims to automate compliance management by leveraging advanced computational techniques Charoenwong et al. (2024). With the emergence of large language models (LLMs), substantial progress has been made in various legal text processing tasks, such as information retrieval, classification, and summarization Guha et al. (2023); Kant et al. (2025). However, directly employing LLMs for compliance checking poses significant challenges. Due to the inherently black-box nature of LLMs and the non-determinism of their outputs, the generated answers may lack consistency and transparency. By contrast, executable code offers clear, verifiable compliance decisions. This raises a fundamental question: Can LLMs bridge the gap between natural language regulations and executable compliance code, thereby enabling fully automated compliance auditing? To explore this, we conducted preliminary experiments with state-of-the-art LLMs, including Gemini-2.5-Pro Kavukcuoglu (2025) and Qwen3-235B Qwen Team (2025). As illustrated in Figure 1, our results indicate that existing LLMs still exhibit significant limitations in regulation-to-code tasks:
  • Incomplete Domain Knowledge: LLMs may conflate domain concepts or overlook essential calculation steps. For instance, Gemini-2.5-Pro detected stock distributions by monitoring overall capital changes, ignoring whether such changes resulted strictly from bonus shares, and miscalculated diluted EPS by skipping post-distribution adjustments, resulting in flawed compliance judgments.
@@
  • Temporal and Logical Inconsistencies: Both models demonstrate limitations in tracking event causality and enforcing regulatory constraints over time. Qwen3-235B defaulted to over-simplified constraint satisfactions in the absence of explicit disclosure data, while Gemini-2.5-Pro wrongly applied event flags across all future periods and relied on ambiguous keyword searches in announcements, failing to uphold event-specific, rule-based compliance.
We attribute these weaknesses primarily to the insufficient domain-specific knowledge of LLMs, and inability to modularize regulatory content, and the absence of appropriate code-oriented training data. Previous research has attempted to address these issues by constructing datasets for improved LLM evaluation in compliance code generation. Existing benchmarks-such as LexGLUE Chalkidis et al. (2022), LegalBench Guha et al. (2023), and MAUD Wang et al. (2023)-focus on tasks like legal text classification, clause extraction, or high-level reasoning. However, they do not modularize financial regulations, and therefore are inadequate for the end-to-end compliance code generation scenario Küçük and Can (2025). Other datasets, such as CODE-ACCORD Hettiarachchi et al. (2025), attempt to formalize condition-action relationships within regulations, but are largely domain-agnostic and not tailored to financial compliance, which frequently relies on specialized terminology. Furthermore, CODE-ACCORD Hettiarachchi et al. (2025) and similar datasets do not effectively reduce model hallucinations in this domain and are predominantly in English, limiting their applicability to diverse regulatory environments. Consequently, there is a lack of a high-quality, Chinese-language dataset for financial compliance checking-a critical resource for advancing LLMs’ understanding, reasoning, and automation capabilities in regulatory contexts.
@@
Experimental Results: We benchmark recent LLMs for regulation structuring and code generation tasks using our dataset’s detailed annotations. DeepSeek-R1-0528 Guo et al. (2025) achieves the best performance on code generation. With Supervised Fine-Tuning (SFT), GLM-4-9B-0414 notably improves its regulation structuring capabilities.
@@ 2.1 RegTech Methods for Financial Compliance
Regulatory Technology (RegTech) leverages computational tools to automate financial compliance Butler and O’Brien (2019). Early RegTech solutions relied on deterministic, rule-based systems and expert models (e.g., GEM Hassan and Logrippo (2009)) using formal logic and standards like XBRL Jeyasingh (2023). While precise and explainable, these symbolic systems struggled with scalability and adaptability amid evolving regulations.
Recent advances in AI, notably Large Language Models (LLMs), have enabled data-driven approaches for parsing regulations and automating compliance workflows Sun et al. (2025b); Jiang et al. (2025). These neural models offer broader generalization but face challenges with domain specificity, rapid regulatory change, and the complexity of legal logic. Key limitations remain in transparency, interpretability, and auditability-essential for compliance contexts Wang et al. (2024a). Even Retrieval-Augmented Generation (RAG) Wiratunga et al. (2024), designed explicitly to adapt to frequent regulatory updates, continues to encounter substantial difficulties due to suboptimal retrieval accuracy and the generator’s limited capacity for complex logical reasoning in compliance checking.
To address this, hybrid pipelines combining neural models with symbolic reasoning or human oversight have become prominent Wang et al. (2024a); Sun et al. (2025a). Our work follows this trajectory, presenting a code-centric approach for LLM-driven compliance automation that enhances rigor, reliability, and legal traceability beyond brittle rule-based pipelines.
@@ 2.2 Related Datasets and Benchmarks
Despite active research in RegTech, there is a critical lack of datasets mapping real regulatory text to executable compliance logic in the financial sector. Most existing resources address subtasks such as document classification (e.g., LEDGAR Tuggener et al. (2020), LexGLUE Chalkidis et al. (2022)), judgment prediction, or QA (e.g., LegalBench Guha et al. (2023), BillSum Kornilova and Eidelman (2019)), but lack annotations for code-level reasoning or logic synthesis.
Datasets closer to financial regulation, like ObliQA Gokhan et al. (2024) and BSARD Lotfi et al. (2025), only support question answering or retrieval. CODE-ACCORD Hettiarachchi et al. (2025) stands out as a logic-annotated resource, but it is limited to building regulations and has narrow domain coverage. Crucially, no existing benchmark, especially in Chinese language systematically couples financial statutes with executable, expert-annotated compliance logic, impeding progress towards true compliance to code. Our work directly addresses this foundational gap.
@@ 3.1 Data Collection and Pre-processing
Compliance-to-Code Dataset is systematically curated from official regulatory documents published by the Beijing Stock Exchange (BSE) [^2] https://www.bse.cn/rule/secnotice_list.html, with a focus on compliance obligations, internal controls, and risk points for publicly listed companies in China. In comparison to the regulations issued by the China Securities Regulatory Commission (CSRC) [^3] http://www.csrc.gov.cn, the BSE provides more detailed and supplementary rules, increasing both the granularity and complexity of regulatory interpretation. We specifically gathered authoritative documents classified as Departmental Rules and Normative Documents, which constitute the regulatory backbone of corporate compliance obligations and reporting requirements. The regulatory corpus reflects all in-force documents as of January 1, 2025, providing a stable but comprehensive snapshot for benchmarking compliance automation.
@@ 3.2 Data Annotation
Drawing inspiration from RegGPT Wang et al. (2024b), we observe that regulations are typically composed of six core elements: meta data (entity/property), constraint, condition, measure, scope, and external references. To facilitate more effective translation of regulatory rules into executable code by LLMs, we propose several refinements to this schema. Specifically, we merge the elements of measure and constraint under the unified category of Constraint, relabel external references as Contextual Information, and rename meta data to Subject.
@@
We reformulated each programming problem as a code completion task and categorized them into three difficulty levels: simple, medium, and difficult. For Simple problems, the LLM needed to fill in one or two missing lines of code; for Medium problems, three lines; and for Difficult problems, the LLM was tasked with generating the entire solution from scratch. DeepSeek-R1 Guo et al. (2025) provided candidate extractions and code suggestions, but all outputs underwent independent expert review and adjudication.
@@ 3.3 Dataset Statistics
Final annotation produced 1,159 CUs (from 1,226 candidates), and 864 directed inter-unit relations - refer to (61.1%), exclude (2.1%), only include (19.7%), should include (17.1%). Each CU has, on average, 0.72 relations, 0.95 subjects, 0.80 conditions, and 0.92 constraints and 0.18 contextual information. (in Figure 2)
These were mapped to 307 distinct executable compliance automation tasks (Python), each code module has a COT reasoning step (average 8 steps/2145 tokens).
@@ 4 FinCheck: A Structured Compliance Checking Pipeline
Figure 3: FinCheck Compliance Checking Pipeline. The framework processes natural language regulations by first using a Structure Predictor to extract key compliance units. These units are then fed into a Code Generator to create verification code. For a specific case, user input triggers an Information Retriever to fetch relevant company data. The generated code is executed with this data, and a Report Generator summarizes the outcome before the final verification result is shown to the user.
@@ 5.2 Baselines
Since there are currently no models specifically designed to convert regulations into code, we are unable to employ any models directly aimed at Task 2. Furthermore, although RegGPT Wang et al. (2024b) demonstrates promising performance for Task 1, it is not open-sourced, which limits our ability to use it in our experiments. As a result, we focus our experiments on general LLMs.
@@ 5.3 Dataset Splits
For our first task, regulation-to-structure, the dataset comprises 1,159 compliance units (CUs), which are divided into training (70%), validation (15%), and test (15%) sets with stratification.
For our second task, structure-to-code, we implemented a strict domain generalization benchmark. We partitioned the available compliance control domains into two disjoint sets. The training set is composed of 249 ’Difficult’ tasks (81.1%) sourced from the first set of domains. The remaining 58 tasks (18.9%) were drawn from the second. To enable a fine-grained evaluation of model performance, we systematically expanded these 58 core tasks into a hierarchical, 307-item test set. This suite, composed of the original 58 ’Difficult’ tasks, plus 34 ’Medium’ and 215 ’Simple’ derived variants, allows for a detailed analysis of how model capabilities vary with complexity.
@@ 5.4 Training Protocols
Base Models: For Task 1, we utilize Qwen3-8B Qwen Team (2025), DeepSeek-R1 and DeepSeek-R1-Distill-Qwen-7B Guo et al. (2025), GLM-4-9B-0414 THUDM (2025), GPT-4.1 OpenAI (2025), Gemini-2.5-pro Kavukcuoglu (2025) and Claude-3.7-Sonnet Anthropic (2025). We also included Fin-R1 Liu et al. (2025), a 7B financial domain model created by fine-tuning Qwen2.5-7B-Instruct on financial reasoning tasks. For Tasks 2 and 3, which focus on code generation, we additionally include Qwen2.5-Coder-7B-Instruct Hui et al. (2024) and DeepSeek-Coder-6.7B-Instruct Guo et al. (2024) in our evaluation.
Supervised Fine-Tuning (SFT): We train using causal language modeling (next-token prediction), AdamW optimizer, cosine learning rate schedule, batch size 1, weight decay 0.01. The learning rate is set at $1\times 10^{-4}$, epochs $1$, top_p $0.75$, and temperature 0. For LoRA, we use rank 8, $\alpha=32$, and dropout $0.05$. We employ Swift  [^5] https://github.com/modelscope/ms-swift for efficiency. The prompt is shown in Appendix H.
@@ 5.5 Evaluation Metrics
For Tasks 2 & 3 (Code Generation), we primarily evaluate Functional Correctness, measured by Pass@1 and Pass@5. Specifically, Pass@k represents the percentage of generated code samples that successfully compile and pass all associated unit tests within the first k attempts. The temperature was set to 0.6, except for DeepSeek-R1-0528, where it was set to 0.0 due to cost considerations. As a secondary metric, we report Code Similarity, quantified by CodeBLEU Ren et al. (2020), which comprehensively assesses the lexical, syntactic, and semantic similarity between generated code and the reference solutions.